# Decoding Dense Embeddings: Sparse Autoencoders for Interpreting and Discretizing Dense Retrieval

### v1

@@ Abstract
Despite their strong performance, Dense Passage Retrieval (DPR) models suffer from a lack of interpretability. In this work, we propose a novel interpretability framework that leverages Sparse Autoencoders (SAEs) to decompose previously uninterpretable dense embeddings from DPR models into distinct, interpretable latent concepts. We generate natural language descriptions for each latent concept, enabling human interpretations of both the dense embeddings and the query-document similarity scores of DPR models. We further introduce Concept-Level Sparse Retrieval (CL-SR), a retrieval framework that directly utilizes the extracted latent concepts as indexing units. CL-SR effectively combines the semantic expressiveness of dense embeddings with the transparency and efficiency of sparse representations. We show that CL-SR achieves high index-space and computational efficiency while maintaining robust performance across vocabulary and semantic mismatches [^1] The code is available at https://github.com/Tro-fish/Decoding-Dense-Embeddings/.
@@ 2.1 Interpretability in Dense Passage Retrieval
Model NMSE MSMARCO Dev TREC DL 2019 TREC DL 2020 MRR@10 Recall@1k NDCG@10 Recall@1k Spearman NDCG@10 Recall@1k Spearman Baseline (SimLM) 0.411 0.986 0.714 0.767 0.697 0.772 Reconstructed (k=32) 0.1903 0.328 0.80/ 0.964 0.98/ 0.589 0.83/ 0.666 0.87/ 0.928 ±$\pm 7.7E\text{-}3$ 0.582 0.84/ 0.681 0.88/ 0.927 ±$\pm 8.1E\text{-}3$ Reconstructed (k=48) 0.1643 0.338 0.83/ 0.968 0.98/ 0.624 0.87/ 0.686 0.89/ 0.936 ±$\pm 6.7E\text{-}3$ 0.613 0.88/ 0.704 0.92/ 0.933 ±$\pm 7.2E\text{-}3$ Reconstructed (k=64) 0.1458 0.347 0.84/ 0.973 0.98/ 0.640 0.90/ 0.692 0.90/ 0.949 ±$\pm 6.2E\text{-}3$ 0.608 0.87/ 0.718 0.93/ 0.948 ±$\pm 6.8E\text{-}3$ Reconstructed (k=128) 0.1069 0.371 0.90/ 0.980 0.99/ 0.664 0.93/ 0.726 0.95/ 0.959 ±$\pm 5.2E\text{-}3$ 0.629 0.90/ 0.734 0.95/ 0.956 ±$\pm 5.4E\text{-}3$
@@ 3.1 Training Sparse Autoencoder
[…]
@@ 3.2 Evaluating Sparse Autoencoder
Additionally, since our goal is to interpret the DPR model via individual latent concepts, we conduct additional experiments to verify whether each latent indeed represents an interpretable semantic concept. Inspired by the “word intrusion test” (Chang et al., 2009), we perform a latent intrusion test, where we collect 9 passages from MSMARCO corpus that most strongly activate a given latent plus 1 randomly chosen “intruder” passage that does not activate the latent. We then use a powerful LLM (GPT-4.1 mini) to identify the outlier. Table 2 shows that, though individual latent’s quality degrades as sparsity decreases, SAEs are still effective at disentangling semantic structures within dense embeddings of DPR models into latent concepts that retain distinct, meaningful information.
@@ 5 Concept-Level Sparse Retrieval
Now we extend the sparse autoencoder framework to the IR domain. We propose Concept-Level Sparse Retrieval (CL-SR), a novel approach that treats each latent concept as a fundamental unit of retrieval. By replacing lexical terms with these semantically coherent and generalized latent concepts, CL-SR enables direct application of term-weighted sparse retrieval algorithms. CL-SR offers Semantic generalization and Computational efficiency. We further analyse these qualities in Section 5.1,  5.2.
Model MSMARCO dev TREC DL 2019 TREC DL 2020 Efficiency MRR@10 Recall@1k NDCG@10 NDCG@10 FLOPs Avg. D Len Storage (GB) Vocab Size (Unsupervised) Sparse Retrieval BM25 0.183 0.853 0.506 0.478 0.13 39.41 0.67 2,660,824 RM3 0.165 0.870 0.522 0.489 39.41 0.67 2,660,824 docT5query 0.277 0.947 0.626 0.607 0.94 756.62 0.98 2,660,824 (Neural) Sparse Retrieval query2doc 0.214 0.918 0.635 0.582 0.87 39.41 0.67 2,660,824 DeepImpact 0.327 0.947 0.657 0.603 71.61 1.40 3,514,102 uniCOIL 0.315 0.924 0.643 0.652 1.03 67.96 1.30 30,522 SPLADEv2 0.340 0.965 0.683 0.671 1.35 91.50 2.60 30,522 CL-SR (Efficient) 0.343 0.954 0.643 0.593 0.11 22.95 0.57 11,709 CL-SR (Max) 0.368 0.969 0.686 0.634 0.74 64.73 1.27 18,679
@@ B.1 Description generation details
Table 6: Document pair statistics for ranking interpretation on TREC DL 2019 and 2020.
Model NMSE TREC DL 2019 TREC DL 2020 NDCG@10 Recall@1k Spearman NDCG@10 Recall@1k Spearman Baseline (TAS-B) 0.721 0.783 0.685 0.800 Reconstructed (k=32) 0.2106 0.573 0.80/ 0.685 0.88/ 0.865 ±$\pm 3.4E-2$ 0.517 0.76/ 0.676 0.85/ 0.861 ±$\pm 3.6E-2$ Reconstructed (k=48) 0.2005 0.575 0.80/ 0.690 0.88/ 0.894 ±$\pm 2.1E-2$ 0.523 0.76/ 0.700 0.88/ 0.891 ±$\pm 2.1E-2$ Reconstructed (k=64) 0.1610 0.593 0.82/ 0.695 0.89/ 0.928 ±$\pm 1.8E-2$ 0.565 0.83/ 0.716 0.90/ 0.925 ±$\pm 1.7E-2$ Reconstructed (k=128) 0.1188 0.626 0.87/ 0.735 0.94/ 0.934 ±$\pm 1.4E-2$ 0.604 0.88/ 0.753 0.94/ 0.932 ±$\pm 1.4E-2$ Baseline (GTR-T5) 0.687 0.737 0.664 0.721 Reconstructed (k=32) 0.1976 0.527 0.77/ 0.617 0.84/ 0.907 ±$\pm 1.4E-2$ 0.469 0.71/ 0.630 0.87/ 0.908 ±$\pm 1.2E-2$ Reconstructed (k=48) 0.1675 0.565 0.82/ 0.631 0.86/ 0.926 ±$\pm 1.1E-2$ 0.547 0.82/ 0.650 0.90/ 0.926 ±$\pm 9.3E-3$ Reconstructed (k=64) 0.1474 0.592 0.86/ 0.647 0.88/ 0.937 ±$\pm 1.0E-2$ 0.568 0.85/ 0.666 0.89/ 0.937 ±$\pm 8.3E-3$ Reconstructed (k=128) 0.1364 0.590 0.86/ 0.687 0.93/ 0.959 ±$\pm 7.3E-3$ 0.614 0.93/ 0.685 0.95/ 0.958 ±$\pm 6.7E-3$
@@
Model TREC-COVID NFCorpus MRR@10 Recall@1k Spearman MRR@10 Recall@1k Spearman Baseline (SimLM) 0.828 0.237 0.498 0.592 Reconstructed (k=32) 0.692 0.83/ 0.188 0.79/ 0.887 ±$\pm 3.3E\text{-}2$ 0.444 0.89/ 0.567 0.96/ 0.806 ±$\pm 2.7E\text{-}2$ Reconstructed (k=48) 0.704 0.85/ 0.190 0.80/ 0.891 ±$\pm 3.5E\text{-}2$ 0.446 0.89/ 0.569 0.96/ 0.809 ±$\pm 2.9E\text{-}2$ Reconstructed (k=64) 0.728 0.88/ 0.193 0.82/ 0.914 ±$\pm 3.2E\text{-}2$ 0.457 0.92/ 0.575 0.97/ 0.851 ±$\pm 6.8E\text{-}2$ Reconstructed (k=128) 0.733 0.89/ 0.216 0.92/ 0.930 ±$\pm 2.8E\text{-}2$ 0.467 0.94/ 0.571 0.96/ 0.896 ±$\pm 2.0E\text{-}2$

### v2

@@ Abstract
Despite their strong performance, Dense Passage Retrieval (DPR) models suffer from a lack of interpretability. In this work, we propose a novel interpretability framework that leverages Sparse Autoencoders (SAEs) to decompose previously uninterpretable dense embeddings from DPR models into distinct, interpretable latent concepts. We generate natural language descriptions for each latent concept, enabling human interpretations of both the dense embeddings and the query-document similarity scores of DPR models. We further introduce Concept-Level Sparse Retrieval (CL-SR), a retrieval framework that directly utilizes the extracted latent concepts as indexing units. CL-SR effectively combines the semantic expressiveness of dense embeddings with the transparency and efficiency of sparse representations. We show that CL-SR achieves high index-space and computational efficiency while maintaining robust performance across vocabulary and semantic mismatches [^1] The code is available at https://github.com/Tro-fish/Decoding-Dense-Embeddings.
@@ 2.1 Interpretability in Dense Passage Retrieval
Model NMSE MSMARCO Dev TREC DL 2019 TREC DL 2020 MRR@10 Recall@1k NDCG@10 Recall@1k Spearman NDCG@10 Recall@1k Spearman Baseline (SimLM) 0.411 0.986 0.714 0.767 0.697 0.772 Reconstructed (k=32) 0.1903 0.328 0.80/ 0.964 0.98/ 0.589 0.83/ 0.666 0.87/ 0.928 $\pm 7.7E\text{-}3$ 0.582 0.84/ 0.681 0.88/ 0.927 $\pm 8.1E\text{-}3$ Reconstructed (k=48) 0.1643 0.338 0.83/ 0.968 0.98/ 0.624 0.87/ 0.686 0.89/ 0.936 $\pm 6.7E\text{-}3$ 0.613 0.88/ 0.704 0.92/ 0.933 $\pm 7.2E\text{-}3$ Reconstructed (k=64) 0.1458 0.347 0.84/ 0.973 0.98/ 0.640 0.90/ 0.692 0.90/ 0.949 $\pm 6.2E\text{-}3$ 0.608 0.87/ 0.718 0.93/ 0.948 $\pm 6.8E\text{-}3$ Reconstructed (k=128) 0.1069 0.371 0.90/ 0.980 0.99/ 0.664 0.93/ 0.726 0.95/ 0.959 $\pm 5.2E\text{-}3$ 0.629 0.90/ 0.734 0.95/ 0.956 $\pm 5.4E\text{-}3$
@@ 3.1 Training Sparse Autoencoder
$\text{Encoder:}\quad z(h)$ $=\sigma(W_{\text{enc}}h+b_{\text{enc}})$ (1)
$\text{Decoder:}\quad\hat{h}(z)$ $=zW_{\text{dec}}+b_{\text{dec}}$ (2)
@@ 3.2 Evaluating Sparse Autoencoder
Additionally, since our goal is to interpret the DPR model via individual latent concepts, we conduct experiments to verify whether each latent indeed represents an interpretable semantic concept. Inspired by the “word intrusion test” (Chang et al., 2009), we perform a latent intrusion test, where we collect 9 passages from MSMARCO corpus that most strongly activate a given latent plus 1 randomly chosen “intruder” passage that does not activate the latent. We then use a powerful LLM (GPT-4.1 mini) to identify the outlier. Table 2 shows that, though individual latent’s quality degrades as sparsity decreases, SAEs are still effective at disentangling semantic structures within dense embeddings of DPR models into latent concepts that retain distinct, meaningful information.
@@ 5 Concept-Level Sparse Retrieval
Now we extend the sparse autoencoder framework to the sparse retrieval domain. We propose Concept-Level Sparse Retrieval (CL-SR), a novel approach that treats each latent concept as a fundamental unit of retrieval. By replacing lexical terms with these semantically coherent and generalized latent concepts, CL-SR enables inverted index retrieval from any dense retrieval model. CL-SR offers Semantic generalization and Computational efficiency. We further analyse these qualities in Section 5.1,  5.2.
Model MSMARCO dev TREC DL 2019 TREC DL 2020 Efficiency MRR@10 Recall@1k NDCG@10 NDCG@10 FLOPs Avg. D Len Storage (GB) Vocab Size (Unsupervised) Sparse Retrieval BM25 0.183 0.853 0.506 0.478 0.13 39.41 0.67 2,660,824 RM3 0.165 0.870 0.522 0.489 39.41 0.67 2,660,824 docT5query 0.277 0.947 0.626 0.607 0.94 756.62 0.98 2,660,824 (Neural) Sparse Retrieval query2doc 0.214 0.918 0.635 0.582 0.87 39.41 0.67 2,660,824 DeepImpact 0.327 0.947 0.657 0.603 71.61 1.40 3,514,102 uniCOIL 0.315 0.924 0.643 0.652 1.03 67.96 1.30 30,522 SPLADE-max 0.340 0.965 0.683 0.671 1.35 91.50 2.60 30,522 CL-SR (Efficient) 0.343 0.954 0.643 0.593 0.11 22.95 0.57 11,709 CL-SR (Max) 0.368 0.969 0.686 0.634 0.74 64.73 1.27 18,679
@@
While more recent variants such as DistilSPLADE and SPLADE v3 (Lassance et al., 2024) achieve stronger performance through heavy distillation and advanced training techniques (e.g., ensemble teacher re-rankers, model-specific hard negative sampling), these enhancements are orthogonal to the focus of our work. To ensure a fair comparison, we report non-distilled SPLADE v2-max as a baseline.
@@ 6 Conclusion
Acknowledgments
This work was partly supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00350379, 30), Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. 2022-0-00369, RS-2022-II220369, 30), Institute of Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2019-II190421, Artificial Intelligence Graduate School Program(Sungkyunkwan University),20), and ICT Creative Consilience Program through the Institute of Information & Communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2020-II201821, 20).
@@ B.1 Description generation details
Table 6: Document pair statistics for ranking interpretation on TREC DL 2019 and $20$.
Model NMSE TREC DL 2019 TREC DL 2020 NDCG@10 Recall@1k Spearman NDCG@10 Recall@1k Spearman Baseline (TAS-B) 0.721 0.783 0.685 0.800 Reconstructed (k=32) 0.2106 0.573 0.80/ 0.685 0.88/ 0.865 $\pm 3.4E-2$ 0.517 0.76/ 0.676 0.85/ 0.861 $\pm 3.6E-2$ Reconstructed (k=48) 0.2005 0.575 0.80/ 0.690 0.88/ 0.894 $\pm 2.1E-2$ 0.523 0.76/ 0.700 0.88/ 0.891 $\pm 2.1E-2$ Reconstructed (k=64) 0.1610 0.593 0.82/ 0.695 0.89/ 0.928 $\pm 1.8E-2$ 0.565 0.83/ 0.716 0.90/ 0.925 $\pm 1.7E-2$ Reconstructed (k=128) 0.1188 0.626 0.87/ 0.735 0.94/ 0.934 $\pm 1.4E-2$ 0.604 0.88/ 0.753 0.94/ 0.932 $\pm 1.4E-2$ Baseline (GTR-T5) 0.687 0.737 0.664 0.721 Reconstructed (k=32) 0.1976 0.527 0.77/ 0.617 0.84/ 0.907 $\pm 1.4E-2$ 0.469 0.71/ 0.630 0.87/ 0.908 $\pm 1.2E-2$ Reconstructed (k=48) 0.1675 0.565 0.82/ 0.631 0.86/ 0.926 $\pm 1.1E-2$ 0.547 0.82/ 0.650 0.90/ 0.926 $\pm 9.3E-3$ Reconstructed (k=64) 0.1474 0.592 0.86/ 0.647 0.88/ 0.937 $\pm 1.0E-2$ 0.568 0.85/ 0.666 0.89/ 0.937 $\pm 8.3E-3$ Reconstructed (k=128) 0.1364 0.590 0.86/ 0.687 0.93/ 0.959 $\pm 7.3E-3$ 0.614 0.93/ 0.685 0.95/ 0.958 $\pm 6.7E-3$
@@
Model TREC-COVID NFCorpus MRR@10 Recall@1k Spearman MRR@10 Recall@1k Spearman Baseline (SimLM) 0.828 0.237 0.498 0.592 Reconstructed (k=32) 0.692 0.83/ 0.188 0.79/ 0.887 $\pm 3.3E\text{-}2$ 0.444 0.89/ 0.567 0.96/ 0.806 $\pm 2.7E\text{-}2$ Reconstructed (k=48) 0.704 0.85/ 0.190 0.80/ 0.891 $\pm 3.5E\text{-}2$ 0.446 0.89/ 0.569 0.96/ 0.809 $\pm 2.9E\text{-}2$ Reconstructed (k=64) 0.728 0.88/ 0.193 0.82/ 0.914 $\pm 3.2E\text{-}2$ 0.457 0.92/ 0.575 0.97/ 0.851 $\pm 6.8E\text{-}2$ Reconstructed (k=128) 0.733 0.89/ 0.216 0.92/ 0.930 $\pm 2.8E\text{-}2$ 0.467 0.94/ 0.571 0.96/ 0.896 $\pm 2.0E\text{-}2$
@@ Appendix F Robustness Test on Mismatch Set
Appendix G Comparison with Neuron-based Interpretability
Previous works like Zhan et al. (2021) has tried to chunk and interpret vector representations in standard basis. We believe this could be problematic since these ranking models are trained with cosine similarity which is rotation invariant. Our method on the other hand does not assume that dimensions have a spatial meaning and instead learns an interpretable basis. Previous works like Olah et al. (2020) also shows that individual neurons (or spatial dimension) are often polysemantic and we believe this holds in our setting as well, motivating our usage of sparse autoencoders. To empirically test this, we performed the latent intrusion test (as described in  4) to the 768-dimensional hidden representation from SimLM. Under the same setting, the accuracy was 0.46, compared to 0.86 with our SAE based method, highlighting a substantial gap in interpretability.
Model Index (GB) Factor Encode Latency (GPU, ms/query) Search Latency (CPU, ms/query) Search Latency (GPU, ms/query) SimLM 26.0 $\times$8.52 3.42 924.42 5.58 SPLADE 2.6 $\times$0.85 3.63 163.71 CL-SR (Max) 1.27 $\times$0.42 3.89 154.25 CL-SR (Efficient) 0.57 $\times$0.19 3.86 117.96
Table 12: Efficiency comparison on MSMARCO Dev. Lower is better. Bold indicates the best in each column.
Appendix H Efficiency Comparison with Dense Retrieval
To provide a more comprehensive comparison, we additionally report an efficiency analysis against the target dense model, SimLM, in Table 12. We provide our experimental results on the MSMARCO Dev set following the implementation of in Li et al. (2022) with PyTorch (GPU, Single Nvidia RTX 3090) and Numpy (CPU, Single Intel(R) Core(TM) i7-14700KF CPU @ 3.40 GHz). We set the batch size to 1 and ran all 6,980 MS MARCO Dev queries three times—reporting the lowest observed latency.
Source Text Top Activated Latents (ID: description (activation)) Query which pathogen depends on living cells 8327: Pathogens causing disease in humans via blood (5.1931) 8442: Immune system influenced by stress and lifestyle (3.3578) 11271: Living as active state or ongoing condition (3.2910) 4166: Definitions of concepts involving multiple components or functions (2.0142) 8190: Response as reaction, answer, or measured outcome (1.9798) 3580: Multifaceted encyclopedic contexts for “Blood” (1.7898) 1322: Defining unicellular organisms by single-cell simplicity (1.6423) Retrieved Incorrect Passage The ability of a multicellular organism to defend itself against invasion by pathogens (bacteria, fungi, viruses, etc.) depends on its ability to mount immune responses. 8306: Definitions of concepts involving multiple components or functions (3.8576) 8442: Immune system influenced by stress and lifestyle (2.9287) 1322: Defining unicellular organisms by single-cell simplicity (2.9263) 8327: Pathogens causing disease in humans via blood (2.8417) 10630: Multifaceted protective actions across legal, psychological, and social domains (1.9795) 6445: Polysemous term "host" across diverse domains (1.9598) 10951: Ability as actualized skill versus potential capacity (1.8653) Not Retrieved Gold Passage Yes, viruses need a living host to replicate. A virion needs to be inside a living cell in order to hijack that cell so that more virus particles can be made by the cell. Since a virus is not a living thing, it doesn’t reproduce in the same way. 10944: Viruses as pseudo-living entities with classification challenges (5.6177) 11271: Living as active state or ongoing condition (4.7194) 8327: Pathogens causing disease in humans via blood (3.6580) 6445: Polysemous term "host" across diverse domains (2.7850) 256: DNA semiconservative replication mechanism (2.3768) 15511: Single-parent reproduction, genetic uniformity, rapid population growth (2.3407) 8296: Immune cells engulfing and digesting foreign material (1.6999)
Table 13: Latent concept activations for a failed query from the MS MARCO Dev set (i.e. R@1000 failure), top-ranked (incorrect) passages, and the gold passage. Each latent concept is listed with its idf-scaled activation value in parentheses.
Appendix I Interpretability Case Study
Table 13 illustrates SAE-based decomposition of SimLM’s dense embeddings for the query “which pathogen depends on living cells,” one of the top-ranked incorrect passage, and the non retrieved gold passage. All three embeddings activate latent 8327 (“pathogens causing disease in humans via blood”), capturing the generic pathogen concept. However, the retrieved (incorrect) passage uniquely shares immune-system latent 8442 (“immune system influenced by stress and lifestyle”) with the query, boosting its matching score to be top-ranked. In contrast, the gold passage activates latent 10944 (“viruses as pseudo-living entities requiring host cells”), which encode the virus-specific host-dependency semantics needed to answer the query; because these latents do not co-occur with the query embedding, SimLM underestimates the passage’s relevance. This case study demonstrates that our SAE-based framework can provide a transparent, concept-level diagnosis of retrieval failures.